If two trains leave New Haven traveling in opposite directions one is traveling at 75 miles per hour and they are 405 miles apart in 3 hours what speed is the second train traveling at.
Found 2 solutions by rfer, stanbon:
Answer by rfer(12644)

405/3=135 apart in one hr
135-75=60 mph

Answer by stanbon(57222)

If two trains leave New Haven traveling in opposite directions one is traveling at 75 miles per hour and they are 405 miles apart in 3 hours what speed is the second train traveling at.
-------------
Let the rate of the 2nd train be "x mph".
---
The separation rate of the two trains is "75+x" mph
------------------------------------------------------
Equation:
3(75+x) = 405
75+x = 135
x = 60 mph (speed of the 2nd train)
=========================================
Cheers,
Stan H.